OpenStudy (anonymous):

When one U.S. Dollar is worth one point four zero Australian Dollars, which of the following could be used as a conversion factor? fifty-sevenths one and two-fifths one and four twenty-fifths two and three-fifths

OpenStudy (anonymous):

one and two fifths 1 + 2/5 = 1 + 0.4 = 1.4